# Extensions of Real-Weighted Fractional Arboricity: Conductance–Resistance Bounds and Monoid Structure

### v4

@@ 6.1 Quantities plotted
For very small graphs ($v\leq 12$ by default) this is computed by brute force enumeration over all connected induced vertex subsets. For larger graphs, the script returns $A_{1}(G)$ only in families where the exact value is known a priori without further optimization: for trees (paths and stars), $A_{1}(G)=1$; for complete graphs, cycles, complete bipartite graphs, and hypercubes, $A_{1}(G)=D_{1}(G)=|E|/(|V|-1)$. The explicit upper bound evaluated at $H=G$ in the unit conductance case comes from Proposition 4.5:
$\mathrm{CSBound}(H{=}G)\;:=\;D_{c}(H=G)\;\leq\;\sqrt{\,\frac{\sum_{e\in E(H)}\dfrac{c(e)}{R_{\text{eff}(e)}^{(G)}}}{\,|V(H)|-1\,}}\,.$
where $R^{(G)}_{\mathrm{eff}}$ is the effective resistance between the endpoints of $e$ in the ambient graph $G$. In the overlay figures we plot the pair $\bigl(A_{1}(G),\mathrm{CSBound}(H{=}G)\bigr)$ against $v=|V|$, and in the tightness figures we plot the ratio $\mathrm{CSBound}(H{=}G)/A_{1}(G)$ (when $A_{1}(G)>0$ and both quantities are finite).
@@ 6.2 Computation on the hypercube family
$R_{d}\;:=\;R_{\mathrm{eff}(e_{0})}^{(Q_{d})}\;=\;R_{\mathrm{eff}(e)}^{(Q_{d})}\qquad\text{for all }e\in E(Q_{d}).$
Hence for an arbitrary conductance assignment $c:E(Q_{d})\to[0,\infty)$,
$\sum_{e\in E(Q_{d})}\frac{c(e)}{R_{\mathrm{eff}(e)}^{(Q_{d})}}\;=\;\frac{1}{R_{d}}\sum_{e\in E(Q_{d})}c(e).$
To compute $R_{\mathrm{eff}(e)}^{(Q_{d})}$, the script follows from [14] and uses a grounded Laplacian solve: relabel vertices as $\{0,1,\dots,|V|-1\}$ with ground node 0, then form the reduced (grounded) Laplacian by deleting the row and column of the Laplacian corresponding to node 0, and for $u,v\in V(G)$, the endpoints of $e$, solve $L_{\mathrm{gr}}x=b$ with $b$ equal to +1+1 at $u$ and 1-1 at $v$ (with the grounded coordinate omitted). The effective resistance is then the resulting voltage drop $R_{\mathrm{eff}(e)}^{(Q_{d})}=|x_{u}-x_{v}|$ [14].
@@ 6.4 Plots
In short, the hypercube family provides a clean stress-test where symmetry allows us to reduce the evaluation of the certificate to a single grounded Laplacian solve (up to relabeling), while still permitting heterogeneous conductance assignments through random sampling. Te overlay figures indicate that $\mathrm{CSBound}(H{=}Q_{d})$ tracks the density functional $D_{c}(Q_{d})$ closely, and the tightness ratios remain near $1$ even as $|V(Q_{d})|=2^{d}$ grows (in particular, the tightness ratio never dips below 1 (Proposition 4.5)). These experiments suggest that the certificate is not merely an artifact of unit weights, but continues to behave stably under significant weight heterogeneity, motivating the broader family-wise computations and complexity discussion in future projects.
@@ 7 Conclusion
together with the associated extremal quantity obtained by maximizing over connected subgraphs. The guiding theme is that embedding conductances into the unweighted arboricity framework produces an invariant that is both structurally similar to fractional arboricity and also naturally tethered to effective resistance and electrical network geometry. On the theoretical side, we developed a usable upper bound for the extremal density in terms of effective resistances in the ambient graph, which is explicit once effective resistances are available. The resulting inequality behaves well under scaling, operates cleanly with classical families, and is computable in practice via standard linear algebra. A brief computational exhibit complements the theory. Using Python, we generated overlay and tightness-ratio plots comparing the baseline density to the explicit certificate across graph sizes. We emphasized the hypercube family $Q_{d}$, where edge–transitivity implies that all edges share the same effective resistance, so $\mathrm{CSBound}(H{=}G)$ can be evaluated from a single grounded–Laplacian solve per dimension. Both in the all-ones base case and under random two-point conductances, the certificate tracks the observed densities closely, and the tightness ratio trends toward $1$ as $d$ grows, supporting the view that the bound is sharp on highly symmetric families.

### v5

@@ 6.1 Quantities plotted
For very small graphs ($v\leq 12$ by default) this is computed by brute force enumeration over all connected induced vertex subsets. For larger graphs, the script returns $A_{1}(G)$ only in families where the exact value is known a priori without further optimization: for trees $T$, including paths and stars, $A_{1}(T)=1$; for complete graphs, cycles, complete bipartite graphs, and hypercubes, $A_{1}(G)=D_{1}(G)=|E|/(|V|-1)$. The explicit upper bound evaluated at $H=G$ in the unit conductance case comes from Proposition 4.5:
$\mathrm{CSBound}(H{=}G)\;:=\;D_{c}(H=G)\;\leq\;\sqrt{\,\frac{\displaystyle\sum_{e\in E(H)}\dfrac{c(e)}{R_{\mathrm{eff}}^{(G)}(e)}}{\,|V(H)|-1\,}}\,.$
where $R_{\mathrm{eff}}^{(G)}(e)$ is the effective resistance between the endpoints of $e$ in the ambient graph $G$. In the overlay figures we plot the pair $\bigl(A_{1}(G),\mathrm{CSBound}(H{=}G)\bigr)$ against $v=|V|$, and in the tightness figures we plot the ratio $\mathrm{CSBound}(H{=}G)/A_{1}(G)$.
@@ 6.2 Computation on the hypercube family
$R_{d}\;:=\;R_{\mathrm{eff}}^{(Q_{d})}(e_{0})\;=\;R_{\mathrm{eff}}^{(Q_{d})}(e)\ \text{for all }e\in E(Q_{d}).$
Hence for an arbitrary conductance assignment $c:E(Q_{d})\to(0,\infty)$,
$\sum_{e\in E(Q_{d})}\frac{c(e)}{R_{\mathrm{eff}}^{(Q_{d})}(e)}\;=\;\frac{1}{R_{d}}\sum_{e\in E(Q_{d})}c(e).$
To compute $R_{\mathrm{eff}}^{(Q_{d})}(e)$, the script follows from [14] and uses a grounded Laplacian solve: relabel vertices as $\{0,1,\dots,|V|-1\}$ with ground node 0, then form the reduced (grounded) Laplacian by deleting the row and column of the Laplacian corresponding to node 0; in other words, let $L_{\mathrm{gr}}$ denote the principal submatrix of $L$ obtained by deleting the row and column corresponding to the ground vertex 0. Then for $u,v\in V(G)$, the endpoints of $e\in E(G)$, solve $L_{\mathrm{gr}}x=b$ with $b$ equal to +1+1 at $u$ and 1-1 at $v$ (with the grounded coordinate omitted). The effective resistance is then the resulting voltage drop $R_{\mathrm{eff}(e)}^{(Q_{d})}=|x_{u}-x_{v}|$ [14].
@@ 6.4 Plots
In short, the hypercube family provides a clean stress-test where symmetry allows us to reduce the evaluation of the certificate to a single grounded Laplacian solve (up to relabeling), while still allowing for heterogeneous conductance assignments through random sampling. The overlay figures indicate that $\mathrm{CSBound}(H{=}Q_{d})$ tracks the density functional $D_{c}(Q_{d})$ closely, and the tightness ratios remain close to $1$ even as $|V(Q_{d})|=2^{d}$ grows (in particular, the tightness ratio never dips below 1 (Proposition 4.5)). These experiments suggest that the certificate is not just an artifact of unit weights, but also behaves stably under significant weight heterogeneity, motivating the broader family-wise computations and complexity discussion in future projects.
@@ 7 Conclusion
together with the associated extremal quantity obtained by maximizing over connected subgraphs. The guiding theme is that embedding conductances into the unweighted fractional arboricity framework produces an invariant that is both structurally similar to fractional arboricity and also naturally tethered to effective resistance in electrical network theory. On the theoretical side, we developed a usable upper bound for the extremal density in terms of effective resistances in the ambient graph, which is explicit once effective resistances are available. The resulting inequality behaves well under scaling, operates cleanly with classical families, and is computable in practice via standard linear algebra. A brief computational exhibit complements the theory. Using Python, we generated overlay and tightness-ratio plots comparing the baseline density to the explicit certificate across graph sizes. We emphasized the hypercube family $Q_{d}$, where edge–transitivity implies that all edges share the same effective resistance, so $\mathrm{CSBound}(H{=}G)$ can be evaluated from a single grounded–Laplacian solve per dimension. Both in the all-ones base case and under random two-point conductances, the certificate is consistently similar in value to the densities, and the tightness ratio trends toward $1$ as $d$ grows, supporting the view that the bound is sharp on highly symmetric families.